# Forge: Foundational Optimization Representations from Graph Embeddings

### v2

@@ 6 Related Work
Recent advances in applying machine learning to combinatorial optimization, particularly Mixed Integer Linear Programming (MILP or MIP), have led to a diverse set of approaches. These methods can be broadly categorized based on their focus areas, such as solver configuration, branching strategies, heuristic design, and generalization across problem types.
@@
Learning-Based Enhancements to MILP Solvers Several works have explored integrating supervised and reinforcement learning into traditional MILP solving pipelines. The survey by Zhang et al. (2023) categorizes these methods into two main groups: those enhancing the Branch-and-Bound process (e.g., branching variable prediction, cutting plane selection, node selection), and those improving heuristic algorithms like Large Neighborhood Search (LNS), Feasibility Pump, and Predict-and-Pick. These models are typically trained end-to-end, with reinforcement learning often relying on imitation learning.
@@
In the context of solver configuration, Hosny & Reda (2024) propose predicting solver parameters by leveraging similarities between problem instances. Their key assumption is that instances with similar costs under one configuration will behave similarly under others. Features include pre-solve statistics and tree-based metrics, with a triplet loss guiding the learning process using solved instance objectives. Kadioglu et al. (2010) propose the ISAC framework that takes a clustering-based approach to instance-specific algorithm configuration. It uses G-means to cluster problem instances and assigns configurations based on cluster membership. The method considers domain-specific features like cost-density ratios and root cost metrics for problems like SCP, MIP, and SAT.
@@
To improve generalization, Boisvert et al. (2024) propose a generic representation for combinatorial problems using abstract syntax trees with 5 node types - variables, constraints, values, operators and a model node. While expressive, this approach results in large, computationally expensive graphs.
@@
Multi-Task and Generalist Models Efforts to unify learning across tasks and problem types have also emerged. Cai et al. (2025a) introduce a multi-task representation learning framework for MIP, training a shared backbone across tasks such as backdoor prediction and solver configuration prediction, followed by fine-tuning for specific problem types. Their method uses a bipartite graph representation, Graph Attention Networks (GAT), and contrastive loss, and is evaluated on problems like CA, MVC and Maximum Independent Set (MIS). Similarly, Drakulic et al. (2024) present GOAL, a generalist agent for combinatorial optimization. It avoids GNNs, instead using mixed attention over edge and node matrices derived from bipartite graphs. Li et al. (2025) propose an LLM-based evolutionary framework that can generate a large set of diverse MILP classes and can be fine tuned to predict integrality gaps and branching nodes.
@@
Graph Neural Networks for Branching and Heuristics Graph-based representations have become standard for encoding MIP instances. One of the earliest work by Gasse et al. (2019) uses GNNs to learn strong branching policies, introducing a bipartite graph structure and dual half-convolutions for message passing between constraints and variables. Chen et al. (2024) revisit GNN for MIPs and show that higher-order GNNs can overcome limitations identified via the 1-Weisfeiler-Lehman test, making all instances tractable for message passing. Cantürk et al. (2024) introduce improvements to the standard GNN workflow for CO so that they generalize on instances of a larger scale than those used in training and propose a two-stage primal heuristic strategy based on uncertainty quantification to automatically configure how solution search relies on the predicted decision values.
@@
Along the lines of Backdoor learning, Cai et al. (2024) use Monte Carlo Tree Search to identify effective backdoors, training a GAT to score variables. Ferber et al. (2022) propose pseudo-backdoors, using one model that characterizes if a subset of variables is a good backdoor and another model to predict whether prioritizing this subset would lead to a smaller run time.
@@
Learning Heuristics and Large Neighborhood Search (LNS) Policies A growing body of work focuses on learning heuristics, particularly for LNS. Huang et al. (2023) use expert hueuristics to create training data followed by random perturbations to create ‘negative’ samples. Then contrastive learning is used to train GATs to predict node probabilities. Other works by Wu et al. (2021) and Song et al. (2020) use deep reinforcement learning to learn destroy operators or decompositions, with rewards based on objective improvements. Khalil et al. (2017) model the success of heuristics at specific nodes by examining instance based characteristics and use logistic regression over a rich feature set, including LP relaxation and scoring metrics. Concretely, they learn a binary classifier to answer whether some heuristic H will find an incumbent solution at a given node N.
@@
Problem Specific Solutions The vehicle routing problem (VRP) has garnered special attention from the community. Zhou et al. (2023) introduce a meta-learning framework for VRPs, enabling generalization across problem sizes and distributions. Berto et al. (2024) explore ML solutions for different kinds of VRPs like those including backhauls, multi-depots, duration limits, mixed backhaul, line hauls etc.. They use a common encoder for all VRP types with global attributes for problem type and local node attributes to capture customer specific attributes like location and demands.
@@
Another problem that has garnered attention is the constraint satisfaction problem. Tönshoff et al. (2023) use GNNs to predict soft assignments, with reinforcement learning rewards based on constraint satisfaction improvements. Duan et al. (2022) propose a contrastive learning framework that generates label-preserving augmentations for SAT problems. These include techniques like unit clause propagation, pure literal elimination, and clause resolution, ensuring that the satisfiability of the instance remains unchanged while enhancing the model’s robustness. Shafi et al. (2025) introduce Graph-SCP, a method that leverages features extracted from both bipartite and hypergraph representations of SCP instances. A GNN is then trained with these features to predict a promising subproblem where the optimal solution is likely to reside. This predicted subproblem is passed to a solver, effectively accelerating the overall solution process.
@@
Unsupervised Approaches Unsupervised learning has also been explored in various forms. Karalias & Loukas (2020) introduce a framework that learns a probability distribution over nodes, optimizing a loss that bounds the probability of finding a solution. These are then decoded using a derandomization process. Bu et al. (2024) build on the work by Karalias & Loukas (2020) by formalizing objective construction and derandomization strategies. They derive explicit formulations tailored to a range of combinatorial problems, including facility location, maximum coverage, and robust graph coloring. Sanokowski et al. (2024) provide an approach for solving combinatorial optimization problems without labeled data by leveraging diffusion models to sample from complex discrete distributions. Their method avoids the need for exact likelihoods by optimizing a loss that upper bounds the reverse KL divergence. While Forge falls in the domain of unsupervised approaches, we differ in that our goal is to not solve a given instance in an unsupervised manner, but rather to learn the graphical structure of various MIP instances in an unsupervised manner followed by supervised fine-tuning to aid in finding the solution.
7 Conclusion & Future Work
We introduce Forge, an unsupervised framework for learning structural representations of MIP instances without requiring access to their solutions. By using a vector-quantized graph autoencoder, Forge effectively captures both local and global structural properties - an essential capability for addressing combinatorial optimization (CO) problems. The discrete codes produced by the vector quantization process behaves as a vocabulary of MIP instances. We demonstrate that these codes can meaningfully cluster previously unseen MIP instances and can be fine-tuned for diverse downstream tasks, including pseudo-cut generation via integrality gap prediction and warm-start variable prediction. Remarkably, a single model with only 3.25 million parameters is used across both tasks and multiple problem types and difficulty levels. We further show that the embeddings produced by Forge can be seamlessly integrated into existing machine learning pipelines for CO, yielding measurable improvements in solver performance. Looking ahead, the instance-level embeddings generated by Forge offer promising opportunities for reinforcement learning-based approaches, such as guiding subproblem selection in Large Neighborhood Search frameworks. Exploring such integrations represents a compelling direction for future work. To support reproducibility and encourage further research, we release all code and pre-trained models.

### v4

@@ A.9 Additional Related Work
Recent advances in applying machine learning to combinatorial optimization, particularly Mixed-Integer Programming (MIP), have led to a wealth of literature. These methods can be broadly categorized based on their focus areas, such as solver configuration, branching strategies, heuristic design, and generalization across problem types.
@@
Learning-based Methods to Enhance MIP Solvers: Several works have explored integrating supervised and reinforcement learning into traditional MIP solving pipelines. The survey by Zhang et al. (2023) categorizes these methods into two main groups: those enhancing the Branch-and-Bound process (e.g., branching variable prediction (Kadioglu et al., 2012; Liberto et al., 2016; He et al., 2014b; Lodi & Zarpellon, 2017; Bengio et al., 2021a), cutting plane selection (Tang et al., 2020; Paulus et al., 2022), node selection He et al. (2014a)), and those improving heuristic algorithms such as Large Neighborhood Search, Feasibility Pump, and Predict-and-Pick. These models are typically trained end-to-end, with reinforcement learning often relying on imitation learning. We refer to (Gast Zepeda et al., 2025) for recent survey.
@@
In the context of solver configuration, the earlier work of  Kadioglu et al. (2010) propose the ISAC framework that takes a clustering-based approach to instance-specific algorithm configuration. It uses g-means to cluster problem instances and assigns configurations based on cluster membership. The method considers domain-specific features such as cost-density ratios and root cost metrics for problems such as SCP, MIP, and SAT. Algorithm selection, scheduling and portfolios have also been studied (Xu et al., 2008; Kadioglu et al., 2011a; Schede et al., 2025; Kemminer et al., 2024). More recently, Hosny & Reda (2024) propose predicting solver parameters by leveraging similarities between problem instances. Their key assumption is that instances with similar costs under one configuration will behave similarly under other configurations. Their features include pre-solve statistics and tree-based metrics, with a triplet loss guiding the learning process using solved instance objectives.
@@
To improve generalization, Boisvert et al. (2024) propose a generic representation for combinatorial problems using abstract syntax trees with five node types - variables, constraints, values, operators and a model node. While expressive, this approach results in large, computationally expensive graphs.
@@
Multi-Task and Generalist Models: Similar to our work here, efforts to unify learning across tasks and problem types have also emerged. Cai et al. (2025a) introduce a multi-task representation learning framework for MIP, training a shared backbone across tasks such as backdoor prediction and solver configuration prediction, followed by fine-tuning for specific problem types. Their method uses a bipartite graph representation, Graph Attention Networks (GAT), and contrastive loss, and is evaluated on problems such as CA, MVC, and (MIS). Similarly, Drakulic et al. (2024) present GOAL, a generalist agent for combinatorial optimization. It avoids GNNs, instead using mixed attention over edge and node matrices derived from bipartite graphs. Li et al. (2025) propose an LLM-based evolutionary framework that can generate a large set of diverse MIP classes and can be fine tuned to predict integrality gaps and branching nodes.
@@
Graph Neural Networks for Branching and Heuristics: Graph-based representations have become standard for encoding MIP instances. One of the earliest work on graph-based learning by Gasse et al. (2019) uses GNNs to learn strong branching policies, introducing a bipartite graph structure and dual half-convolutions for message passing between constraints and variables. Chen et al. (2024) revisit GNN for MIPs and show that higher-order GNNs can overcome limitations identified via the 1-Weisfeiler-Lehman test, making all instances tractable for message passing. Cantürk et al. (2024) introduce improvements to the standard GNN workflow for CO so that they generalize on instances of a larger scale than those used in training and propose a two-stage primal heuristic strategy based on uncertainty quantification to automatically configure how solution search relies on the predicted decision values.
@@
Along the lines of Backdoor learning, Cai et al. (2024a) use Monte Carlo Tree Search to identify effective backdoors, training a GAT to score variables. Ferber et al. (2022) propose pseudo-backdoors, using one model that characterizes if a subset of variables is a good backdoor and another model to predict whether prioritizing this subset would lead to a smaller run time.
@@
Learning Heuristics and Large Neighborhood Search (LNS): Earlier works such as (Kadioglu et al., 2017; 2011b) proposes learning reactive restart and impact-based strategies to improve search. Recently, a growing body of work focuses on learning heuristics, particularly for Large Neighborhood Search (LNS). Huang et al. (2023) use expert heuristics to create training data followed by random perturbations to create ‘negative’ samples. Then contrastive learning is used to train GATs to predict node probabilities. Other works by Wu et al. (2021) and Song et al. (2020) use deep reinforcement learning to learn destroy operators or decompositions, with rewards based on objective improvements. Khalil et al. (2017) model the success of heuristics at specific nodes by examining instance based characteristics and use logistic regression over a rich feature set, including LP relaxation and scoring metrics. Cai et al. (2025b) propose Balans, an adaptive meta-solver for MIPs with online learning capability that does not require any supervision or apriori training. During the search, the selection among different neighborhood definitions is guided on the fly via multi-armed bandit algorithms (Strong et al., 2021; 2019).Yilmaz et al. (2025) extend Balans Cai et al. (2024b) using solver- and algorithmic-level parallelism into ParBalans to improve performance on challenging MIP instances.
@@
Problem Specific Methods: The vehicle routing problem (VRP) has garnered special attention from the community (Berto et al., 2025; Hottung & Tierney, 2022).Zhou et al. (2023) introduce a meta-learning framework for VRPs, enabling generalization across problem sizes and distributions. Berto et al. (2024) explore ML solutions for different kinds of VRPs like those including backhauls, multi-depots, duration limits, mixed backhaul, line hauls, among others. They use a common encoder for all VRP types with global attributes for problem type and local node attributes to capture customer specific attributes such as location and demands.
@@
Another problem that has garnered attention is the constraint satisfaction problem. Tönshoff et al. (2023) use GNNs to predict soft assignments, with reinforcement learning rewards based on constraint satisfaction improvements. Duan et al. (2022) propose a contrastive learning framework that generates label-preserving augmentations for SAT problems. These include techniques such as unit clause propagation, pure literal elimination, and clause resolution, ensuring that the satisfiability of the instance remains unchanged while enhancing the model’s robustness. Shafi et al. (2025) introduce Graph-SCP, a method that leverages features extracted from both bipartite and hypergraph representations of SCP instances. A GNN is then trained with these features to predict a promising subproblem where the optimal solution is likely to reside. This predicted subproblem is passed to a solver, effectively accelerating the overall solution process.
@@
Unsupervised Approaches: Unsupervised learning has also been explored in various forms. Karalias & Loukas (2020) introduce a framework that learns a probability distribution over nodes, optimizing a loss that bounds the probability of finding a solution. These are then decoded using a derandomization process. Bu et al. (2024) build on the work by Karalias & Loukas (2020) by formalizing objective construction and derandomization strategies. They derive explicit formulations tailored to a range of combinatorial problems, including facility location, maximum coverage, and robust graph coloring. Sanokowski et al. (2024) provide an approach for solving combinatorial optimization problems without labeled data by leveraging diffusion models to sample from complex discrete distributions. Their method avoids the need for exact likelihoods by optimizing a loss that upper bounds the reverse KL divergence. While Forge falls in the domain of unsupervised approaches, we differ in that our goal is to not solve a given instance in an unsupervised manner, but rather to learn the graphical structure of various MIP instances in an unsupervised manner followed by supervised fine-tuning to aid in finding the solution.